how far from the base of the building must a 26 foot ladder be placed so that it reaches 10 feet up the wall
Answer by user_dude2008(1862)

x^2+10^2=26^2
x^2+100=676
x^2=676-100
x^2=576
x=sqrt(576)
x=24 <------------- Answer